# Simulating Nighttime Visible Satellite Imagery of Tropical Cyclones Using Conditional Generative Adversarial Networks

### v2

@@ II-B Band selection
Figure 2: Weighting functions of AHI Band08–10 and Band16 under U.S. Standard Atmosphere and a satellite zenith angle of 0°.
@@
Satellite viewing angles also create differences in the correspondence of VIS and IR images, especially for the full-disk data. The satellite viewing angles affect the atmospheric absorption and the projected lengths of lights and shades at the edge of the full-disk observation. Additionally, the satellite viewing angles contribute to the approximate simulation of the glare of the sun. Therefore, we included the satellite zenith angle and the satellite azimuth angle in our input datasets. We also included a basemap extracted from the red channel of NASA’s Blue Marble Next Generation [39] which provides information about land usually not being available in IR bands.
@@ III-A The revised pix2pix model
Compared to the original pix2pix model, our model is designed for more complex application scenarios. For input $X$, instead of having $256\times 256$ input images with 3 channels, we have $512\times 512$ input images with 12 channels. If the original U-net structure is applied, the receptive field will be too small compared to the size of input $X$. Therefore, To expand the receptive field and capture global features more comprehensively, we adopted a deeper U-Net architecture. Additionally, we increased the number of dropout layers to mitigate overfitting, especially given the complexity of the input data.
@@
We changed the value of the coefficient to $k=20$ in this study so that the weight of the second term in Eq. (10) remains approximately the same as the one in Eq. (5). A flow chart of our revised pix2pix model is shown in Fig. 5.
Figure 5: Flow chart of our revised pix2pix model.
@@ III-C Model validation and statistical comparison
Since our model is eventually used for simulating nighttime visible imagery, we tested the model performance during nighttime using both AHI data and VIIRS DNB data. This approach of using DNB imagery as a substitute for visible imagery at nighttime is also adopted by [16] and [18].
@@
The presence of the glare of the moon affects model validation. Since our model uses the satellite viewing angles of AHI data, it cannot simulate the glares observed by VIIRS. We calculated the angular distance between the center of the image and where the glare would be the most intense as follows:
@@ IV-A Daytime validation performance
U-Net [^2] The U-Net and U-net++ models have relatively inconsistent performance at nighttime according to [12]. 0.85 - 0.09 - - -
U-Net+++++ + [^2] The U-Net and U-net++ models have relatively inconsistent performance at nighttime according to [12]. 0.86 - 0.07 - - -
@@
Figure 9: Sample daytime image pairs and their corresponding difference images and 2D histograms. The color scale in the third column matches Fig. 1.
Fig. 9 shows 5 AI-generated images in the validation set, their paired real VIS images, the difference images between AI-generated and VIS images, and the corresponding 2D histograms. The 5 sample images are Typhoon Nepartak (2016) at 23:10 UTC on July 5, Typhoon Trami (2018) at 04:20 UTC on September 22, Typhoon Yutu (2018) at 23:00 UTC on October 23, Typhoon Haishen (2020) at 23:00 UTC on September 3, and Cyclone Lucas (2021) at 07:40 UTC on February 6 in sequence. This covers a variety of different solar zenith angles, solar azimuth angles, seasons, latitudes, and TC intensities. The images of Cyclone Lucas demonstrate that our model is capable of simulating land regions with certain reflectivity. The 2D histograms show that most of the pixels are concentrated around the $y=x$ line, indicating an outstanding overall model performance. We also noticed that the 2D histograms show very few pixels with AI-generated reflectance above 0.95, while the real VIS images do have reflectance $\sim$1 at some pixels. This might have caused the slight negative Bias at low solar zenith angles as shown in Table IV.
@@ IV-B Nighttime validation performance
Figure 13: Sample nighttime image pairs and their corresponding difference images and 2D histograms. The color scale in the third column matches Fig. 1.
Fig. 13 presents AI-generated imagery and normalized DNB imagery in pairs, their difference images, and the 2D histograms. To minimize the parallax between satellites, images near the equator were chosen. Since most TCs are away from the equator, no TCs were included. The sample images are at 15:18 UTC on February 28, 2018, 15:10 UTC on April 28, 2018, 14:29 UTC on May 27, 2018, 14:36 UTC on August 25, and 14:12 UTC on September 22, 2018 in sequence. The image pairs have a RMSE varying from 0.036 to 0.071 depending on the complexity of cloud features. The 2D histograms show that most pixels align closely around the $y=x$ line, showing a clear correspondence between AI-generated and normalized DNB images.
The nighttime statistical result in Table VI is slightly worse than the daytime statistical result in Table III, which we mainly attribute to the minor parallax at the edge of the image caused by the rapidly varying viewing angles of VIIRS. Fig. 14 zooms in on the difference image from the fifth row of Fig. 13, in which several regions affected by the parallax are identified. The mismatch of images caused by parallax creates streak-like patterns on the difference image, contributing a considerable part of the RMSE between the two images. While the parallax negatively impacts the statistical results, the cloud features remain accurately simulated, albeit in slightly different locations. Additionally, although the geographical location of clouds simulated using AHI IR data mismatches that of VIIRS data, it is still consistent with AHI VIS data. Therefore, the existence of parallax in cross-satellite nighttime validation does not affect the accuracy and utility of AI-generated images.
Figure 14: The difference image from the fifth row and third column of Fig. 13, with three regions affected by parallax circled.
In conclusion, despite the statistical results being slightly worse at nighttime, the AI-generated images remain highly accurate, demonstrating successful model generalization from daytime to nighttime. We attribute this success to the relatively low diurnal temperature variation in tropical ocean regions, which supports the fundamental assumption that the IR bands have consistent $T_{b}$ from daytime and nighttime.
@@ V Discussion
This study presents a CGAN model that converts IR data, direction parameters of the sun and the satellite, and land basemap into highly accurate AI-generated VIS reflectance. As a result, the CGAN model significantly enhances nighttime TC monitoring when VIS images are unavailable. Currently, we provide real-time AI-generated images of TCs to the public (https://ai-vis.dapiya.top).
@@
Beyond basic cloud visualizations, we aim to expand our model’s applications by using AI-generated reflectance to derive Atmospheric Motion Vectors (AMVs). AMVs are estimations of winds derived from successive geostationary satellite images [58]. The derivation of low-level AMVs relies on VIS images to track the low-level clouds accurately. When applied to model TC surface wind field, low-level AMVs derived from VIS images show lower RMSE, lower Bias, and improved coverage compared to the AMVs derived from IR images [59]. With the capability to identify low clouds as VIS images do, our AI-generated images could be assimilated as satellite data to complement the derivation of nighttime low-level AMVs. An important application of low-level AMVs is to model the size of the TC wind field. With the AI-generated images as input, the TC wind field can be modeled accurately and continuously at night. Further studies are needed to confirm the feasibility of this potential application.

### v3

@@ II-B Band selection
Figure 2: Weighting functions of AHI Band08–10 and Band16 under Standard Tropical Atmosphere and a satellite zenith angle of 0°.
@@
Satellite viewing angles also create differences in the correspondence of VIS and IR images, especially for the full-disk data. The satellite viewing angles affect the atmospheric absorption and the projected lengths of lights and shades at the edge of the full-disk observation. Additionally, the satellite viewing angles contribute to the approximate simulation of the sun glare. Therefore, we included the satellite zenith angle and the satellite azimuth angle in our input datasets. We also included a basemap extracted from the red channel of NASA’s Blue Marble Next Generation [39] which provides information about land usually not being available in IR bands.
@@ III-A The revised pix2pix model
Figure 5: Detailed architecture of the revised pix2pix model. Note that the VIS images are replicated to 3 channels, enabling future framework generalization for multi-color imagery.
Compared to the original pix2pix model, our model is designed for more complex application scenarios. Therefore, we modified the architectures of both G and D.
For input $X$ of G, instead of having $256\times 256$ images with 3 channels, we have $512\times 512$ images with 12 channels. If the original U-Net structure were applied directly, the receptive field would be too small relative to the size of $X$. To expand the receptive field and more comprehensively capture global features, we adopted a deeper U-Net architecture with nine downsampling layers. For activation functions, we use LeakyReLU in the encoder to extract subtle features from the input, ReLU in the first eight layers of the decoder for efficient reconstruction, and Tanh in the final decoding layer for normalization. Additionally, we increased the usage of dropout to mitigate overfitting, especially given the complexity of the IR data. Specifically, dropout is introduced in eleven deep layers of U-Net.
The input of D concatenates $X$ with either $Y$ or $Y^{\prime}$, forming $512\times 512$ images with 15 channels. Note that Y is replicated to 3 channels, which allows framework generalization for multi-color imagery (e.g. true-color imagery). With a similar consideration of receptive fields, we use a deeper network with six convolutional layers, each employing LeakyReLU as the activation function. Fig. 5 illustrates the details of the revised model architecture.
@@
We changed the value of the coefficient to $k=20$ in this study so that the weight of the second term in Eq. (10) remains approximately the same as the one in Eq. (5).
@@ III-C Model validation and statistical comparison
Since our model is eventually used for simulating nighttime visible imagery, we tested the model performance during nighttime using both AHI data and VIIRS DNB data. This approach of using DNB imagery as a substitute for visible imagery at night is also adopted by [16] and [18].
@@
The presence of the moon glare affects model validation. Since our model uses the satellite viewing angles of AHI data, it cannot simulate the glares observed by VIIRS. We calculated the angular distance between the center of the image and where the glare would be the most intense as follows:
@@ IV Results
U-Net [^2] The U-Net and U-net++ models have relatively inconsistent performance at night according to [12]. 0.85 - 0.09 - - -
U-Net+++++ + [^2] The U-Net and U-net++ models have relatively inconsistent performance at night according to [12]. 0.86 - 0.07 - - -
@@
IV-A Daytime validation performance
Figure 9: Sample daytime image pairs and their corresponding difference images and 2D histograms. The color scale in the third column matches Fig. 1, and the dashed lines in the fourth column are the best-fit lines of linear regression.
Fig. 9 shows 5 AI-generated images in the validation set, their paired real VIS images, the difference images between AI-generated and VIS images, and the corresponding 2D histograms. The 5 sample images are Typhoon Nepartak (2016) at 23:10 UTC on July 5, Typhoon Trami (2018) at 04:20 UTC on September 22, Typhoon Yutu (2018) at 23:00 UTC on October 23, Typhoon Haishen (2020) at 23:00 UTC on September 3, and Cyclone Lucas (2021) at 07:40 UTC on February 6 in sequence. This covers a variety of different solar zenith angles, solar azimuth angles, seasons, latitudes, and TC intensities. The images of Cyclone Lucas demonstrate that our model is capable of simulating land regions with certain reflectivity. We performed linear regression for each of the image pairs and overlaid the best-fit lines on the 2D histograms. In all 5 cases, the best-fit lines lie very close to the $y=x$ line, with $R^{2}$ values exceeding 0.9995, indicating an outstanding and consistent model performance. We also noticed that the 2D histograms show very few pixels with AI-generated reflectance above 0.95, while the real VIS images do have reflectance $\sim$1 at some pixels. This might have caused the slight negative Bias at low solar zenith angles as shown in Table IV.
@@ IV-B Nighttime validation performance
Fig. 14 presents AI-generated imagery and normalized DNB imagery in pairs, their difference images, and the 2D histograms. To minimize the parallax between satellites, images near the equator were chosen. Since most TCs are away from the equator, no TCs were included. The sample images are at 15:18 UTC on February 28, 2018, 15:10 UTC on April 28, 2018, 14:29 UTC on May 27, 2018, 14:36 UTC on August 25, and 14:12 UTC on September 22, 2018 in sequence. The image pairs have a RMSE varying from 0.036 to 0.071 depending on the complexity of cloud features. The linear regressions yield $R^{2}$ values above 0.9944 for all 5 image pairs, showing a strong correspondence between AI-generated and normalized DNB images. The best-fit lines have slopes slightly below 1, likely because the residual moon glare slightly increases the measured reflectance of the clear sea surface.
@@
Figure 13: The sample difference image at 14:12 UTC on September 22, 2018, with three regions affected by parallax circled.
Figure 14: Sample nighttime image pairs and their corresponding difference images and 2D histograms. The color scale in the third column matches Fig. 1, and the dashed lines in the fourth column are the best-fit lines of linear regression.
The nighttime statistical result in Table VI is slightly worse than the daytime statistical result in Table III. Besides the minor effect of moon glare, we primarily attribute this decline to parallax at the image edges, which is caused by VIIRS’s rapidly varying viewing angles. Fig. 13 zooms in on the difference image from the fifth row of Fig. 14, in which several regions affected by the parallax are identified. The mismatch of images caused by parallax creates streak-like patterns on the difference image, contributing a considerable part of the RMSE between the two images. While the parallax negatively impacts the statistical results, the cloud features remain accurately simulated, albeit in slightly different locations. Additionally, although the geographical location of clouds simulated using AHI IR data mismatches that of VIIRS data, it is still consistent with AHI VIS data. Therefore, the existence of parallax in cross-satellite nighttime validation does not affect the accuracy and utility of AI-generated images.
In conclusion, despite the statistical results being slightly worse at night, the AI-generated images remain highly accurate, demonstrating successful model generalization from daytime to nighttime. We attribute this success to the relatively low diurnal temperature variation in tropical ocean regions, which supports the fundamental assumption that the IR bands have consistent $T_{b}$ from daytime and nighttime.
@@ V Discussion
This study presents a CGAN model that converts IR data, direction parameters of the sun and the satellite, and land basemap into highly accurate AI-generated VIS reflectance. As a result, the CGAN model significantly enhances nighttime TC monitoring when VIS images are unavailable. Currently, we provide real-time AI-generated images of TCs to the public (https://ai-vis.dapiya.top). The complete source code and datasets used in this study are publicly available at our GitHub repository (https://github.com/Dapiya/AI-VIS).
@@
Beyond pixel-level simulations, we aim to incorporate broader meteorological metrics to enhance model validation and explore diverse model applications. Unlike current metrics such as SSIM and PSNR, object-based verification metrics can directly identify cloud patches and evaluate their similarity in real and AI-generated imagery from multiple aspects. For example, the Method for Object-based Diagnostic Evaluation (MODE) [58] can be adapted for cloud reflectance and evaluate cloud simulations based on shape, position, and reflection intensity.
VIS imagery is widely used to derive meteorological information, including cloud optical depth and atmospheric motion vectors (AMVs). VIS reflectance plays a key role in retrieving cloud optical depth, which subsequently enables the derivation of cloud phase and cloud-top properties [59, 60]. The performance of our model can be further validated by substituting AI-generated reflectance into cloud property retrieval algorithms and comparing the results with those obtained using real VIS reflectance. Additionally, since cloud property retrieval is less accurate at night due to the absence of reflective bands, AI-generated reflectance could be combined with recent methods for simulating MWIR bands [61] to enhance nighttime retrieval accuracy.
AMVs are estimations of winds derived from successive geostationary satellite images [62]. The derivation of low-level AMVs relies on VIS images to track the low-level clouds accurately. When applied to model TC surface wind field, low-level AMVs derived from VIS images show lower RMSE, lower Bias, and improved coverage compared to the AMVs derived from IR images [63]. With the capability to identify low clouds as VIS images do, our AI-generated images could be assimilated as satellite data to complement the derivation of nighttime low-level AMVs. An important application of low-level AMVs is to model the size of the TC wind field. With the AI-generated images as input, the TC wind field can be modeled accurately and continuously at night. Further studies are needed to confirm the feasibility of this potential application.